# SoK: Evaluating Jailbreak Guardrails for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have achieved remarkable progress, but their deployment has exposed critical vulnerabilities, particularly to jailbreak attacks that circumvent safety mechanisms. Guardrails—external defense mechanisms that monitor and control LLM interactions—have emerged as a promising solution. However, the current landscape of LLM guardrails is fragmented, lacking a unified taxonomy and comprehensive evaluation framework. In this Systematization of Knowledge (SoK) paper, we present the first holistic analysis of jailbreak guardrails for LLMs. We propose a novel, multi-dimensional taxonomy that categorizes guardrails along six key dimensions, and introduce a Security-Efficiency-Utility evaluation framework to assess their practical effectiveness. Through extensive analysis and experiments, we identify the strengths and limitations of existing guardrail approaches, explore their universality across attack types, and provide insights into optimizing defense combinations. Our work offers a structured foundation for future research and development, aiming to guide the principled advancement and deployment of robust LLM guardrails.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of applications, revolutionizing fields from natural language understanding to content generation [1, 2, 3, 4, 5, 6, 7, 8]. However, their increasing sophistication and widespread adoption have also exposed significant vulnerabilities. A prominent concern is their susceptibility to jailbreak attacks [9, 10], where adversaries craft malicious inputs to bypass safety alignments and elicit harmful, biased, or unintended responses. The proliferation of such attacks underscores the urgent need for robust defense mechanisms. Among various defense strategies, guardrails [11, 12, 13] have emerged as a promising approach, aiming to monitor and control LLM interactions without altering the underlying model’s weights or core functionalities.
Guardrail-based defenses offer a distinct advantage over prompt-based [14] or tuning-based [15, 16] methods as they can effectively filter jailbreak attempts while preserving the integrity of the target LLM’s original output capabilities. Despite their potential, the current landscape of LLM guardrails is characterized by siloed innovation. Numerous research teams and organizations have proposed various guardrail solutions, often tailored to specific scenarios, attack vectors (e.g., focusing primarily on single-turn attacks), or proprietary systems. This ad-hoc development has resulted in a fragmented ecosystem of defense mechanisms, lacking a unified understanding or a systematic classification framework to position and compare these disparate efforts.
@@
To address these critical gaps, this Systematization of Knowledge (SoK) paper provides the first comprehensive analysis and structuring of the rapidly evolving field of jailbreak guardrails for LLMs. We aim to consolidate the disparate research efforts, offering a clear and structured understanding of the current state-of-the-art. Our primary contributions are threefold: (1) we propose a novel, multi-dimensional taxonomy for classifying LLM guardrails, enabling a nuanced understanding of their design characteristics; (2) we introduce a holistic evaluation framework centered on the Security-Efficiency-Utility trifecta, promoting more practical and comprehensive assessments; and (3) we conduct extensive analysis based on our framework, yielding valuable insights into the performance of existing guardrails and identifying promising avenues for future research.
Specifically, our contributions are as follows:
@@
    • Security Granularity: Defining the scope of the guardrail detection (Token-level, Sequence-level, or Session-level).
@@
    • Interpretability: Assessing the transparency of the guardrail’s decision-making process and noting that increased interpretability might inadvertently introduce new attack surfaces or reasoning errors.
  • A Security-Efficiency-Utility Evaluation Framework: We introduce a novel framework for evaluating guardrails that balances three crucial aspects:
@@
    • We conduct a tri-objective (Security-Efficiency-Utility) evaluation of mainstream guardrail methods to identify balanced solutions and those effective against diverse jailbreak categories.
    • We investigate specific hypotheses, such as the efficacy of session-level guardrails against multi-turn attacks, the influence of intervention stage on latency, the impact of technical paradigms on resource consumption, and the relationship between security granularity and utility.
@@ 2 Jailbreak Attacks in LLMs
[…]
$\sup_{P\in\mathcal{T}^{\star}}\Pr_{R\sim\text{LLM}(P)}[\text{JUDGE}(R,G)=\text{True}]$ (1)
@@
Existing Jailbreak Attacks. Jailbreak attacks can be broadly categorized into two types: single-turn and multi-turn jailbreaks. Single-turn jailbreaks involve crafting a single prompt to elicit harmful responses, while multi-turn jailbreaks exploit the interactive nature of LLMs by engaging in a dialogue with the model over multiple turns. Due to the maturity of research on single-turn attacks and the relative scarcity of multi-turn attack studies, we further divide single-turn attacks into four types: manual methods, optimization-based approaches, generation-based strategies, and implicit jailbreaks.
@@
  • Optimization-based Jailbreaks. These methods iteratively refine adversarial prompts using techniques like gradient-based optimization or search strategies [9, 22, 23, 10, 24]. GCG [9] introduced a greedy coordinate gradient method to optimize adversarial suffixes, enabling transferable jailbreaks across models and prompts. Sitawarin et al. [22] extended this with GCG++, leveraging a proxy model to enhance optimization. Beyond gradient-based techniques, JSAA [23] employed random search for suffix optimization, while AutoDAN [10] used a hierarchical genetic algorithm to create human-readable jailbreak prompts. RLbreaker [25] utilized reinforcement learning to efficiently search for adversarial prompts, outperforming stochastic methods like JSAA and AutoDAN.
@@
  • Implicit Jailbreaks. These techniques disguise malicious intent within query text to bypass LLM safety mechanisms [33, 34, 35, 36, 37, 38, 39, 18, 40]. For instance, Handa et al. [33] demonstrated word substitution as a simple evasion method. DrAttack [34] decomposes harmful prompts into smaller, less detectable sub-prompts. Puzzler [35] embeds clues within queries to guide the LLM toward producing harmful outputs indirectly. Another approach involves translating harmful prompts into languages where LLM safety mechanisms are weaker [36, 37, 38, 39, 18, 40]. Deng et al. [36] and Yong et al. [37] found that low-resource languages, such as Zulu, often exhibit less robust safety alignment. Obfuscation techniques, including encoding or encrypting harmful prompts, further reduce LLM sensitivity to malicious inputs [41, 18, 40].
  • Multi-turn Jailbreaks. One multi-turn attack strategy is the fine-grained task decomposition, which decomposes the original malicious query into several less harmful sub-questions [42, 43, 44]. While this decomposition strategy successfully circumvents current safety mechanisms, it may be easily mitigated by including these finer-grained harmful queries in safety training data. Alternatively, researchers propose to use human red teamers to expose vulnerabilities of LLMs against multi-turn attacks [45]. Moreover, Yang et al. [46] depends on the heuristics from [27] and its seed examples to implement its attacks. Crescendo [47] gradually steers benign initial queries towards more harmful topics. The implementation of Crescendo is based on the fixed and human-crafted seed instances, making it challenging to generate diverse and effective attacks. By contrast, ActorAttack [48] proposes to discover diverse attack clues inside the model’s prior knowledge. X-Teaming [49] achieves more effective and diverse multi-turn attacks by adaptive collaborative agents for planning, attack optimization, and verification.
@@ 3.1 Jailbreak Guardrail Definition
Alternatively, if the guardrail inspects the model’s output $R=\text{LLM}(P)$, a successful jailbreak is characterized by:
@@
This highlights that a successful adversary must not only craft a prompt that bypasses the LLM’s internal safety measures but also deceives the guardrail into permitting the harmful interaction or content.
@@ 3.3 Guardrail Evaluation Framework
  • Attack Success Rate (ASR): ASR measures the proportion of adversarial attempts that successfully bypass the guardrail and elicit harmful or unintended responses from the target LLM. Formally, it is defined as the percentage of attack queries for which the LLM system equipped with the guardrail fails to block or mitigate the attack. A lower ASR indicates stronger defense.
@@
    Extra $\text{Extra Delay}=T_{\text{guardrail + LLM}}-T_{\text{LLM}}$ (4)
@@
    GPU $\text{GPU Overhead}=M_{\text{guardrail + LLM}}-M_{\text{LLM}}$ (5)
@@ 4.1 Intervention Stages
On one hand, intra-processing guardrails rely on gradient information to identify potential jailbreak attempts. These methods analyze the gradients of the model’s inputs or parameters during inference to identify unusual patterns or anomalies that may indicate adversarial inputs. For example, GradSafe [58] computes the similarity between the input’s gradient w.r.t. the safety-critical parameters and the unsafe reference gradients. Gradient Cuff [61] compare the gradient norm of refusal loss w.r.t. the query prompt with a threshold. Token Highlighter [80] uses the gradient norm of the affirmation loss for each token in the user query to locate the jailbreak-critical tokens.
On the other hand, intra-processing guardrails can analyze the model’s internal states for jailbreak detection [65, 66, 67, 75, 76, 85]. These methods leverage the model’s hidden states, or other internal representations to identify patterns indicative of jailbreak attempts. For example, Circuit Breaking [67] interrupts the LLM to output harmful content when harmful states are detected. JBShield [85] analyzes the differences of the LLM’s internal states between the jailbreak prompts and the benign queries. These approaches can provide more nuanced insights into the model’s behavior and vulnerabilities, enabling more effective detection of sophisticated jailbreak techniques. However, these approaches typically require white-box access to the model, which limits their applicability to open-source LLMs or scenarios where model internals are accessible.
@@
In addition to assess the semantic harmfulness of the response, some methods cleverly use the difference in responses caused by destroying the adversarial properties of the jailbreak prompt as the basis for judgment [57, 59, 88, 89]. For example, SmoothLLM [57] randomly perturbs/permutes multiple copies of a given input prompt to generate a set of responses, and then vote by the perturbed responses to determine the safety of the original request. SemanticSmooth [59] and JailGuard [89] perform more complex mutations instead of simple character-level changes like SmoothLLM, such as paraphrasing or translation into other languages. MirrorShield [88] generates mirror prompts that preserve the syntactic structure of the input while ensuring semantic safety. Beyond directly assessing the semantic harmfulness of the response, an alternative category of methods ingeniously leverages the discrepancies in outputs that arise from disrupting the adversarial characteristics of the initial jailbreak prompt [57, 59, 88, 89]. SmoothLLM [57], for instance, operates by randomly perturbing or permuting multiple copies of a given input prompt to generate a set of responses; the safety of the original request is then determined by a voting mechanism based on these perturbed responses. Advancing this concept, SemanticSmooth [59] and JailGuard [89] implement more complex mutations than the simple character-level alterations used by SmoothLLM, such as paraphrasing the prompt or translating it into other languages. In a similar vein, MirrorShield [88] generates ”mirror” prompts that aim to preserve the syntactic structure of the input while ensuring its semantic safety. These mutation-based methods capitalize on the inherent properties of jailbreak prompts to identify adversarial inputs, rendering them particularly effective against sophisticated attacks. Nevertheless, they may incur additional computational overhead due to the requirement for numerous response evaluations or intricate input transformations.

### v2

@@ Abstract
Large Language Models (LLMs) have achieved remarkable progress, but their deployment has exposed critical vulnerabilities, particularly to jailbreak attacks that circumvent safety alignments. Guardrails—external defense mechanisms that monitor and control LLM interactions—have emerged as a promising solution. However, the current landscape of LLM guardrails is fragmented, lacking a unified taxonomy and comprehensive evaluation framework. In this Systematization of Knowledge (SoK) paper, we present the first holistic analysis of jailbreak guardrails for LLMs. We propose a novel, multi-dimensional taxonomy that categorizes guardrails along six key dimensions, and introduce a Security-Efficiency-Utility evaluation framework to assess their practical effectiveness. Through extensive analysis and experiments, we identify the strengths and limitations of existing guardrail approaches, provide insights into optimizing their defense mechanisms, and explore their universality across attack types. Our work offers a structured foundation for future research and development, aiming to guide the principled advancement and deployment of robust LLM guardrails.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of applications, revolutionizing fields from natural language understanding to content generation [1, 2, 3, 4, 5, 6]. However, their increasing sophistication and widespread adoption have also exposed significant vulnerabilities. A prominent concern is their susceptibility to jailbreak attacks [7, 8], where adversaries craft malicious inputs to bypass safety alignments and elicit harmful, biased, or unethical responses. The proliferation of such attacks underscores the urgent need for robust defense mechanisms. Among various defense strategies, guardrails [9, 10, 11] have emerged as a promising approach, aiming to monitor and control LLM interactions without altering the underlying model’s weights or core functionalities.
Guardrail-based defenses offer a distinct advantage over other defense methods (e.g., tuning-based approaches [12]) as they can effectively filter jailbreak attempts while preserving the integrity of the target LLM’s original output capabilities. Despite their potential, the current landscape of LLM guardrails is characterized by siloed innovation. Numerous research teams and organizations have proposed various guardrail solutions, often tailored to specific scenarios, attack vectors (e.g., focusing primarily on single-turn attacks), or proprietary systems. This ad-hoc development has resulted in a fragmented ecosystem of defense mechanisms, lacking a unified understanding or a systematic classification framework to position and compare these disparate efforts.
@@
To address these critical gaps, this Systematization of Knowledge (SoK) paper provides the first comprehensive analysis and structuring of the rapidly evolving field of jailbreak guardrails for LLMs. We aim to consolidate the disparate research efforts, offering a clear and structured understanding of the current state-of-the-art. Our primary contributions are threefold: (1) we propose a novel, multi-dimensional taxonomy for classifying LLM guardrails, enabling a nuanced understanding of their design characteristics; (2) we introduce a holistic evaluation framework centered on the Security-Efficiency-Utility trifecta, promoting more practical and comprehensive assessments; and (3) we conduct extensive analysis based on our framework, yielding valuable insights into the performance of existing guardrails and identifying promising avenues for future research. Specifically, our contributions are as follows:
@@
    • Safety Granularity: Defining the scope of the guardrail detection (Token-level, Sequence-level, or Session-level).
@@
    • Interpretability: Assessing the transparency of the guardrail’s decision-making process.
  • A Security-Efficiency-Utility (SEU) Evaluation Framework: We introduce a novel framework for evaluating guardrails that balances three crucial aspects:
@@
    • We conduct a tri-objective (SEU) evaluation of mainstream guardrail methods to identify balanced solutions and those effective against diverse jailbreak categories.
    • We investigate specific hypotheses, such as the efficacy of session-level guardrails against multi-turn attacks, the influence of intervention stage on latency, the impact of technical paradigms on resource consumption, and the relationship between safety granularity and utility.
@@ 2 Jailbreak Attacks in LLMs
Jailbreak Formulation. The jailbreak phenomenon indicates that specific malicious instructions can bypass the safety mechanisms of LLMs, leading to the generation of harmful or unethical outputs. This is particularly concerning as it highlights the potential for adversaries to exploit vulnerabilities in LLMs to produce toxic or harmful content. The jailbreak process can be viewed as a two-step procedure: (1) crafting an adversarial prompt $P$ that elicits a harmful response from the LLM, and (2) evaluating the generated response $R$ against a predefined harmful objective $G$ using a classifier JUDGE. JUDGE returns ‘True’ if the generated response $R$ meets the harmful objective $G$, i.e., $\text{JUDGE}=\text{True}$, otherwise ‘False’. Let $\mathcal{T}$ represent the LLM’s vocabulary. Formally, we can define the classifier $\text{JUDGE}:\mathcal{T}^{\star}\times\mathcal{T}^{\star}\to\{\text{True},\text{False}\}$. The adversary’s goal is to maximize the probability of generating responses classified as satisfying the harmful objective $G$. This can be expressed mathematically as:
$\sup_{P\in\mathcal{T}^{\star}}\Pr_{R\sim\text{LLM}(P)}[\text{JUDGE}(R,G)=\text{True}],$ (1)
@@
Existing Jailbreak Attacks. Jailbreak attacks can be broadly categorized into two types: single-turn and multi-turn jailbreaks. Single-turn jailbreaks involve crafting a single prompt to elicit harmful responses, while multi-turn jailbreaks exploit the interactive nature of LLMs by engaging in a dialogue with the model over multiple turns. Due to the maturity of research on single-turn attacks and the relative scarcity of multi-turn attack studies, a further breakdown of the single-turn attack is necessary. Building on prior works [13, 11] that categorize jailbreak attacks by their technical paradigms, we divide single-turn attacks into similar four types: manual, optimization-based, generation-based and implicit jailbreaks.
@@
  • Optimization-based Jailbreaks. These methods iteratively refine adversarial prompts using techniques like gradient-based optimization or search strategies [7, 19, 20, 8, 21]. GCG [7] introduced a greedy coordinate gradient method to optimize adversarial suffixes, enabling transferable jailbreaks across models and prompts. Sitawarin et al. [19] extended this with GCG++, leveraging a proxy model to enhance optimization. Beyond gradient-based techniques, JSAA [20] employed random search for suffix optimization and BON [22] induces jailbreaks by repeatedly sampling augmentations of a harmful instruction until one succeeds, while AutoDAN [8] used a hierarchical genetic algorithm to create human-readable jailbreak prompts. RLbreaker [23] utilized reinforcement learning to efficiently search for adversarial prompts, outperforming stochastic methods like JSAA and AutoDAN.
@@
  • Implicit Jailbreaks. This class of attacks is derived from the Linguistics-based and Encoding-based jailbreak categories identified in JBShield [13]. Given that both strategies conceal malicious intent within the query text to circumvent LLM safety mechanisms, they are collectively categorized as implicit jailbreaks. For instance, Handa et al. [31] demonstrated word substitution as a simple evasion method. DrAttack [32] decomposes harmful prompts into smaller, less detectable sub-prompts. Puzzler [33] embeds clues within queries to guide the LLM toward producing harmful outputs indirectly. Another approach involves translating harmful prompts into languages where LLM safety mechanisms are weaker [34, 35, 36, 37, 15, 38]. Deng et al. [34] and Yong et al. [35] found that low-resource languages, such as Zulu, often exhibit less robust safety alignment. Obfuscation techniques [39], including encoding or encrypting harmful prompts, further reduce LLM sensitivity to malicious inputs [15, 38].
  • Multi-turn Jailbreaks. One multi-turn attack strategy is the fine-grained task decomposition, which decomposes the original malicious query into several less harmful sub-questions [40, 41, 42]. While this decomposition strategy successfully circumvents current safety mechanisms, it may be easily mitigated by including these finer-grained harmful queries in safety training data. Alternatively, researchers propose to use human red teamers to expose vulnerabilities of LLMs against multi-turn attacks [43]. Moreover, Yang et al. [44] depend on the heuristics from [25] and its seed examples to implement their attacks. Crescendo [45] gradually steers benign initial queries towards more harmful topics. The implementation of Crescendo is based on the fixed and human-crafted seed instances, making it challenging to generate diverse and effective attacks. By contrast, ActorAttack [46] proposes to discover diverse attack clues inside the model’s prior knowledge. X-Teaming [47] achieves more effective and diverse multi-turn attacks by adaptive collaborative agents for planning, attack optimization, and verification.
@@ 3 Definition, Taxonomy & Evaluation
3.1 Jailbreak Defense
Jailbreak defense comprises methods to protect LLMs from jailbreak attacks intended to elicit harmful or policy-violating content. We classify these defenses based on their point of application into four categories: prompt-based, tuning-based, refining-based defenses, and guardrails, which respectively focus on modifying prompts, optimizing model parameters, controlling model outputs, and external monitoring.
Prompt-based defenses alter the input prompt to enhance the LLM’s adherence to safety guidelines [48, 16, 49, 50]. For instance, Self-Reminder [48] prepends ethical reminders, while ICD [16] provides examples of harmful queries and safe responses. Tuning-based defenses optimize the LLM’s parameters to improve its inherent safety. This includes techniques like reinforcement learning from human feedback (RLHF) [51, 52], machine unlearning to erase harmful knowledge [53], model editing to patch vulnerabilities [54], supervised fine-tuning on curated safety datasets [55, 56], and adversarial training to bolster robustness against attacks [12]. Refining-based defenses steer the output generation process in real-time [57, 58, 59]. For example, RAIN [57] uses token-level evaluation to guide rewinds for safer generation, whereas Aligner [59] employs a lightweight model to correct initial responses towards alignment. Guardrails are external modules that monitor and control the interaction [9, 10, 60]. As illustrated in Figure 1, they can operate on inputs (pre-processing), internal model states (intra-processing), or outputs (post-processing). Distinct from the other categories, guardrails do not modify the target LLM’s inference; rather, they detect harmfulness and filter the interaction to prevent jailbreaks.
@@ 3.2 Jailbreak Guardrail Definition
Alternatively, if the guardrail inspects the model’s internal features $F$ or output $R=\text{LLM}(P)$, a successful jailbreak is characterized respectively by:
$\text{Assess}(\mathcal{G}_{R},F)=\textit{allow}~\land~\text{JUDGE}(R,G)=\text{True},$ (3)
@@
This highlights that a successful adversary must not only craft a prompt that bypasses the LLM’s internal safety measures but also deceive the guardrail into permitting the harmful interaction or content.
@@ 3.4 Guardrail Evaluation Framework
  • Attack Success Rate (ASR): ASR measures the proportion of adversarial attempts that successfully bypass the guardrail and elicit harmful or unethical responses from the target LLM. Formally, it is defined as the percentage of attack queries for which the LLM system equipped with the guardrail fails to block or mitigate the attack. A lower ASR indicates a stronger defense.
@@
    $\text{Extra Delay}=T_{\text{guardrail + LLM}}-T_{\text{LLM}},$ (5)
@@
    $\text{GPU Overhead}=M_{\text{guardrail + LLM}}-M_{\text{LLM}},$ (6)
@@
Alignment with Taxonomy Dimensions. The proposed taxonomy in §3.3 is intentionally structured to align with the empirical evaluation dimensions of the SEU framework, which form the core contribution of this work. Particularly, we would analyze the first three taxonomy dimensions (Intervention Stage, Technical Paradigm, and Safety Granularity) in depth, because they have a direct, measurable impact on Security, Efficiency, and Utility, forming the foundation of our quantitative analysis. In contrast, the latter three dimensions (i.e., Reactiveness, Applicability, and Explainability) are more qualitative in nature and represent critical future research directions for enhancing the adaptability, deployment flexibility, and transparency of guardrail systems. By focusing our empirical analysis on the dimensions most directly tied to operational performance, we provide a principled and actionable foundation for evaluating and optimizing jailbreak guardrails in practice.
@@ 4.1 Intervention Stages
On one hand, intra-processing guardrails rely on gradient information to identify potential jailbreak attempts. These methods analyze the gradients of the model’s inputs or parameters during inference to identify unusual patterns or anomalies that may indicate adversarial inputs. For example, GradSafe [68] computes the similarity between the input’s gradient w.r.t. the safety-critical parameters and the unsafe reference gradients. Gradient Cuff [71] compares the gradient norm of refusal loss w.r.t. the query prompt with a threshold. Token Highlighter [90] uses the gradient norm of the affirmation loss for each token in the user query to locate the jailbreak-critical tokens.
On the other hand, intra-processing guardrails can analyze the model’s internal states for jailbreak detection [75, 76, 77, 85, 86, 13]. These methods leverage the model’s hidden states or other internal representations, to identify patterns indicative of jailbreak attempts. For example, Circuit Breaking [77] interrupts the LLM to output harmful content when harmful states are detected. JBShield [13] analyzes the differences of the LLM’s internal states between the jailbreak prompts and the benign queries. These approaches can provide more nuanced insights into the model’s behavior and vulnerabilities, enabling more effective detection of sophisticated jailbreak techniques. However, these approaches typically require white-box access to the target model, which limits their applicability to open-source LLMs or scenarios where model internals are accessible.
@@
Beyond directly assessing the semantic harmfulness of the response, an alternative category of methods ingeniously leverages the discrepancies in outputs that arise from disrupting the adversarial characteristics of the initial jailbreak prompt [67, 69, 97, 98]. SmoothLLM [67], for instance, operates by randomly perturbing or permuting multiple copies of a given input prompt to generate a set of responses; the safety of the original request is then determined by a voting mechanism based on these perturbed responses. Advancing this concept, SemanticSmooth [69] and JailGuard [98] implement more complex mutations than the simple character-level alterations used by SmoothLLM, such as paraphrasing the prompt or translating it into other languages. In a similar vein, MirrorShield [97] generates “mirror” prompts that aim to preserve the syntactic structure of the input while ensuring its semantic safety. These mutation-based methods capitalize on the inherent properties of jailbreak prompts to identify adversarial inputs, rendering them particularly effective against sophisticated attacks. Nevertheless, they may incur additional computational overhead due to the requirement for numerous response evaluations or intricate input transformations.
@@
RQ 1: Pre-processing guardrails can reject harmful inputs before they reach the target LLM, intra-processing mechanisms operate concurrently during LLM inference, and post-processing techniques must await LLM’s outputs. This raises a pertinent question: To what extent does the specific intervention stage of a guardrail, be it pre-processing, intra-processing, or post-processing, influence overall response latency?